Suppose you are given a 10.7 mM solution of lactose. You then hydrolyze the sample (i.e. break up the lactose, disaccharide, into two monosaccharides). Assuming the efficiency of this hydrolysis was 60% (i.e. only 60% of the lactose molecules were hydrolyzed), what is the difference in reducing sugar concentration (i.e. as measured by Somogyi-Nelson test) between the starting lactose solution and the final hydrolyzed solution? Assume that there is a minimum change in total volume through this process. Give you answer in mM to 3 significant digits.
Answer #1

Lactose <===> Glucose + Galactose

Intial 10.7mM. 0 0

Change -c c c

Equilibrium c(1-) c c

Now total concernatation = c(1-)+ 2c = c + c​​​​​​

= 0.6, c = 10.7 mM, total concernatation = 10.7+ 10.7 × 0.6 =  17.120 mM